OpenStudy (anonymous):

Write the formula for the geometric sequence. 5, 10, 20, 40, …

hartnn (hartnn):

\(A_n = A_1r^{n-1}\)

OpenStudy (anonymous):

Try to find the common ratio first by b/a or c/b or d/c

hartnn (hartnn):

where A1 is first term, r = common ratio

hartnn (hartnn):

could u find common ratio ?

OpenStudy (anonymous):

so is the formula A1=5*2^n?

hartnn (hartnn):

its n-1 in formula

OpenStudy (anonymous):

ohhhh ok so A1=5*2^n-1

hartnn (hartnn):

yes, An = 5* 2^(n-1)

hartnn (hartnn):

or An= 5/2 * 2^n

OpenStudy (anonymous):

ok c:

hartnn (hartnn):

thats your formula